# RAGServe: Fast Quality-Aware RAG Systems with Configuration Adaptation

### v1

@@ 6 Implementation
We implement RAGServe in about 1.5K lines of code in Python. We build RAGServe on top of the state-of-the-art popular LLM serving engine vLLM [38]. We also use supporting modules from Langchain [8] in order to have efficient implementations of the multiple synthesis methods. For RAGServe’ LLM used for configuration space pruning, we use OpenAI’s API [54] to invoke GPT-4o on the RAG queries and HuggingaceAPI [71] to use LLama-3.1-70B as models to profile the queries.
@@
We use a state-of-the-art embedding method Cohere embed-v3.0 [4] We perform FAISS [16] L2-distance similarity search on the embeddings of the chunks relative to the query, in order to retrieve relevant chunks for RAG inference.
Finally, we use PyTorch’s [5] library modules support to perform query-level memory profiling and measurement to implement the best-fit scheduling logic and request batching.
@@
Table 1: Input and output length (# of tokens) distributions of the RAG datasets used in our evaluation.
7 Evaluation
The key takeaways from the evaluation are
  • $\bullet$
    Lower delay : Across 4 task representative datasets for RAG QA, RAGServe achieves 1.642.54×1.64-2.54$\times$1.64 - 2.54 × lower response delay compared to fixed configurations of comparable quality.
  • $\bullet$
    Higher throughput : RAGServe achieves 1.84.5×1.8-4.5$\times$1.8 - 4.5 × higher throughput than RAG serving systems which use fixed configurations reaching similar quality.
  • $\bullet$
    Negligible overhead : RAGServe’ profiler’s delay is negligible compared to the overall delay of the LLM’s RAG inference.
Figure 10: RAGServe achieves 1.642.54×1.64-2.54$\times$1.64 - 2.54 × lower delay compared to both best fixed configuration baselines and quality-optimized RAG configuration without sacrificing generation quality.
Figure 11: RAGServe achieves 1.84.5×1.8-4.5$\times$1.8 - 4.5 × higher throughput (at 1.8 seconds) than baselines which use fixed configurations of closest (not higher) quality.
7.1 Setup
Models and hardware: : We evaluate RAGServe on a popular model for LLM inference, specifically the fine-tuned version of Mistral-7B-v3. We also use Llama3.1-70B for additional experiments. All models are fine-tuned such that they can take long contexts (up to 32K and 128K respectively). We apply AWQ-model quantization both models. We use an NVIDIA A40 GPU server with 2 GPUs to benchmark our results. The server is equipped with 384GB of memory and two Intel(R) Xeon(R) Gold 6130 CPUs with Hyper-threading and Turbo Boost enabled by default. We use 1 GPU to serve Mistral-7B-v3 and 2 GPUs to serve Llama3.1-70B.
Datasets:  We use multiple RAG QA datasets which have a variety of different query profiles, in order to have task-representative workloads. Table 1 summarizes their input-output length statistics.
  • $\bullet$
    Squad [59]: Squad is a reading comprehension dataset, consisting of questions on Wikipedia articles, where the answer to every question is a segment of text, or span, from the corresponding reading passage. Squad is classified as a single-hop reasoning dataset.
  • $\bullet$
    Musique [68]: Musique is a multihop QA dataset with reasoning-based questions. It is designated to test LLM’s multi-hop reasoning ability where one reasoning step critically relies on information from another.
  • $\bullet$
    KG RAG FinSec [45]: KG RAG Finsec is part of a Knowledge Graph family of RAG datasets and focuses on financial domain questions from Fortune 500 companies. This dataset contains quarterly financial reports and queries need to read information for multiple chunks for answering.
  • $\bullet$
    QMSUM [82]: QMSUM is a human-annotated benchmark for query-based multi-domain meeting summarization that is designed to test LLM’s reasoning-based summarization capabilities. This dataset consists of multiple meeting transcripts and the queries are to summarize relevant spans of meetings from them.
For our experiments, we build a retrieval database from all of the datasets. We split the queries’ contexts into fixed-sized chunks using Langchain [8] for the database. We use a state-of-the-art embedding method Cohere embed-v3.0 [4] We perform FAISS [16] L2-distance similarity search on the embeddings of the chunks relative to the query, in order to retrieve relevant chunks for RAG inference.
To simulate a real RAG workload, we create a mix of queries from each dataset, and send the queries to RAGServe using arrival rates that follow a Poisson distribution. We report the results per dataset.
@@ 7.1 Setup
  • F1-score is used to evaluate the RAGServe’s serving model’s response in the datasets. It measures the probability that the generated answer matches the ground-truth answer of the question-answering task. It is the most widely adopted metric for evaluating RAG QA tasks [61, 63, 10]
@@
  • Delay is used to measure the generation response delay of the model for every RAG query. We choose this metric similar to other RAG serving papers [40, 66, 62]
  • Dollar Cost is used to measure the lower cost of using RAGServe’s profiler as compared to using larger serving models with fixed configurations having the closest accuracy.
Baselines:  We compare RAGServe with the following baselines.
  • vLLM: Every RAG query is served with the same fixed RAG configuration with vLLM and we compare with several such configurations.
  • AdaptiveRAG*: We implement AdaptiveRAG’s [31], query complexity-based RAG-configuration choosing, and choose the configuration which maximizes the F1-score.
@@
  • Parrot*: We implement Parrot’s [40] configuration-based batching. Parrot* does not adapt the configuration per query. We compare with Parrot* using fixed RAG configurations which achieve the closest quality to us.
7.2 Overall improvement
@@
Lower delay without sacrificing generation quality:  Figure 10 shows RAGServe achieves delay reduction 1.642.54×1.64-2.54$\times$1.64 - 2.54 × over AdaptiveRAG* with no reduction in F1-score. Over using fixed configurations of similar delay, served with both Parrot* and vLLM, RAGServe achieves $12-18\%$ higher F1-score.
@@
Higher throughput at lower delay:  Figure 11 shows RAGServe achieves higher throughput compared to fixed-configuration baselines when they choose the fixed-config which achieves the closest quality. Compared to Parrot* and vLLM, RAGServe achieves 1.84.5×1.8-4.5$\times$1.8 - 4.5 × times higher throughput.
@@
Understanding RAGServe’ improvement:  RAGServe’s gains come from being able to jointly select the configuration based on the available resource, along with perform batching and scheduling used by existing RAG serving systems. RAGServe achieves higher quality than the fixed-config baselines as it is able to adapt the RAG-configuration per query, focusing on quality. At the same time, it reduces delay by scheduling per-input configurations better than fixed configurations which achieve similar quality.
@@
RAGServe achieves higher throughput due to being able to adapt configurations based on resource availability as compared to the baselines. Both Parrot* and vLLM schedule fixed RAG-configurations and cannot benefit from delay achieved by adapting the configuration like RAGServe. Parrot* can improve the delay over using fixed configurations with vLLM by 1.41.8×1.4-1.8$\times$1.4 - 1.8 × but cannot improve the quality.
7.3 Analyzing the gains from RAGServe
@@
Delay saving:  Figure 12 shows the contribution of every component of RAGServe. We compare with vLLM plus a fixed configuration, which achieves the highest quality (orange bar). If we just use the profiler’s ranges and naively the fixed median value every time (red bar), we achieve 1.41.68×1.4-1.68$\times$1.4 - 1.68 × reduction in delay. Further, we see the effect of only batching (like Parrot*), by choosing the median value configuration and batching, we achieve 1.11.2×1.1-1.2$\times$1.1 - 1.2 × reduction in delay. Finally, RAGServe achieves even greater delay reduction by 1.451.75×1.45-1.75$\times$1.45 - 1.75 × by adapting the configuration based on available GPU memory along with batching.
Figure 12: Understanding the reasons for delay improvement in RAGServe.
@@
Figure 13: Even with increasing the inference model size, fixed configurations have 2.386.8×2.38-6.8$\times$2.38 - 6.8 × higher cost and lower quality compared to RAGServe.
@@
Figure 14: Improvement for RAGServe using feedback from the output helps improve the F1-score by $4-6\%$.
@@
Figure 15: RAGServe achieves lower delay by 2.12.4×2.1-2.4$\times$2.1 - 2.4 × at the same quality even with a larger inference LLM.
@@
Figure 16: Breakdown analysis: By tuning more knobs in RAGServe, we can see better quality-delay tradeoffs.
@@
Figure 17: RAGServe’ performance gains remain substantial even with a smaller LLM profiler.
@@
Figure 18: RAGServe’ profiler delay is at most 1/10th of end-to-end response delay across queries from all datasets.
@@
Cost saving:  Figure 13 shows RAGServe (including its profiler) has significant lower dollar cost saving and higher F1-score, compared to choosing the best fixed configuration, with increasing model complexity. The cost of using a (LLama3-70B) inference model with vLLM and a fixed configuration is higher by 2.38×2.38$\times$2.38 × times while also having a lower F1-score of $6.5\%$ times across datasets. Even more powerful models like GPT-4o fail to achieve the same F1-score with fixed configurations but have a much higher cost of 6.8×6.8$\times$6.8 ×.
@@
Profiler feedback-based improvement:  In Figure 14 we show the effect of the golden-configuration-based feedback to the profiler in order to improve its output. We use a 350 query sample for the QMSUM and KG RAG FinSec dataset as the workload. We see that with the feedback mechanism (blue line), the F1-score improves by $4-6\%$ as compared to not having feedback (red line) from the outputs of the golden configuration. It is important to note that if the feedback mechanism cannot result in the output of very expensive configurations, as RAGServe’ joint scheduler will ensure it doesn’t pick increasingly expensive configurations based on the GPU resource constraint.
7.4 Sensitivity analysis
@@
Changing the inference LLM:  Figure 15 shows the outcome of changing the inference LLM to a larger LLM (Llama3.1-70B) on the Musique and QMSUM datasets. Even with a more powerful LLM, RAGServe achieves 2.12.4×2.1-2.4$\times$2.1 - 2.4 × lower delay than AdaptiveRAG* at a similar F1-score. The best fixed-configuration baselines such as Parrot* and vLLM have a lower F1-score of $7-10\%$. In RAG, models mainly rely on the external context to answer the question instead of the model weights and we only get a $2\%$ improvement in F1-score compared to the smaller inference models.
@@
Incrementally using knobs in RAGServe:  In Figure 16, we show the benefit we the improvement we get by incrementally adding more knobs to RAGServe. We measure this for the QMSUM dataset with the original Mistral-7B-v3 model. We first only tune the num_chunks (red point). Progressively we tune the RAG-configuration knobs of synthesis_method and intermediate_length and scheduling. We achieve $5,4,3\%$ higher F1-Score compared to vLLM. Finally, by adding the scheduling, 2.8×2.8$\times$2.8 × lower delay reduction in delay.
@@
Changing the profiler LLM:  Figure 17 shows the effect of changing the LLM profiler from GPT-4o to a less powerful, Llama3.1-70B model. RAGServe with the new profiler, still achieves 1.42.1×1.4-2.1$\times$1.4 - 2.1 × over AdaptiveRAG* with a similar F1-score. Compared to fixed configurations with Parrot* and vLLM which achieve similar delay, RAGServe achieves $10-14\%$ higher F1-score.
Changing the embedding algorithm:  RAGServe uses for retrieval on quality. RAGServe picks a state-of-art retrieval algorithm Cohere embed-v3.0 [4]. When compare with two more popular retrieval algorithms All-mpnet-base-v2 from SentenceBERT [60] and text-embedding-3-large-256 from OpenAI [18], the F1-score difference is less than 1%, when we use RAGServe’ configuration selection. The delay has no measurable difference as the retrieval is >100×>100$\times$> 100 × faster than LLM synthesis [6].
@@
Delay overhead of RAGServe’s per-query profiling:  We show the negligible delay overhead of using an LLM profiler with RAGServe to serve RAG queries Figure 18 shows the fraction of RAGServe’ profiler of the total end-to-end delay. Using the profiler at most adds less $0.1$ fraction of the total delay and in the average case, the profiler only adds $0.03-0.06$ fraction across queries from all datasets.
8 Related work
@@
Systems for serving RAG:  Several systems have been proposed for RAG [40, 34, 66, 48, 37, 31, 77, 17, 2, 32, 79] which focus on improving retrieval using complex, iterative retrieval algorithms or on serving model selection. RAGServe can work in conjunction with such systems as RAGServe focuses on optimizing quality and serving latency, independent of how the retrieval algorithm identifies chunks for retrieval.
@@
KV cache storage and retrieval:  Storing and reusing KV cache across different requests have been commonly studied in recent work [44, 28, 41, 14, 43, 56, 2, 22, 76, 65, 81, 38]. RAGServe can work alongside these systems, where instead of retrieving chunks, it can retrieve the KV Caches for generating the output. In RAG, some additional optimizations are needed to combine KV Caches of different chunks that don’t share a common prefix. This is important as the trivial concatenation of KV Caches loses important cross-attention and reasoning between chunks. These optimizations are enabled by KV Cache blending-based approaches [75, 29, 9, 35, 70, 26]. However RAG workloads have a large number of related contexts across queries and storing all the KV Cache is extremely expensive. We do not measure the KC Cache reuse ratio across queries and leave it for future work.
@@
Prefill-Decode Optimizations:  Several systems have proposed optimizations to speed-up prefill and decode for LLMs by leveraging unique properties of each phase [3, 83, 72, 58, 33, 11, 64, 84]. Some notable techniques include chunked-prefill which allows interleaving prefill and decode requests and disaggregated prefill which separates compute nodes for prefill and decode. All of these optimizations enable faster generation speed but don’t focus on generation quality. RAGServe is orthogonal to these LLM serving systems optimizations.
9 Limitations
RAGServe is designed to work with commonly deployed RAG pipelines. New research directions in RAG [78, 17] have shown developed more complex pipelines and configuration knobs for deep chain-of-thought RAG workloads. While their pipelines improve these on complex workloads, they achieve similar performance on the commonly used RAG workloads we consider [1]. Additionally, some RAG systems use external re-rankers [23, 46] to choose the exact num_chunks the model might need. However, for joint scheduling, it is still necessary to have the opportunity to tune this number and RAGServe can be adapted to use a range of chunks from the re-ranker output.
10 Conclusion

### v3

@@ 6. Implementation
We implement METIS in about 2K lines of code in Python on top of the state-of-the-art popular LLM serving engine vLLM (kwon2023efficient, ). For the profiler used for configuration space pruning, we define a class LLMProfiler inheriting OpenAI’s Chat Completion API (OpenAIAPI, ) interface (to invoke GPT-4o) and HuggingaceAPI (Wolf_Transformers_State-of-the-Art_Natural_2020, ) (to invoke LLama-3.1-70B) as models to profile the queries.
@@
We use Cohere-embed-v3.0 (cohere_embed, ) as a state-of-the-art embedding method. We construct a FAISS (douze2024faiss, ) index using the IndexFlatL2 interface and perform L2-distance similarity search with index.search(query_embedding, top_k) on the chunk embeddings to retrieve for RAG inference. We use the LLMChain interface from Langchain (Chase_LangChain_2022, ) in order to build efficient implementations of multiple synthesis methods.
Finally, we use PyTorch’s (Ansel_PyTorch_2_Faster_2024, ) library modules support to perform query-level memory profiling and measurement to implement the best-fit scheduling logic and request batching. Particularly, we use pynvml to construct get_free_memory() with its interfaces of nvmlDeviceGetHandleByIndex and nvmlDeviceGetMemoryInfo to measure the amount of GPU memory available. We measure the current num-seqs and num-batched-tokens within vLLM to calculate which configuration can be fit into the current batch, based on the GPU availability and the request’s memory requirement.
7. Evaluation
The key takeaways from the evaluation are
  • $\bullet$
    Lower delay : Across 4 task representative datasets for RAG QA, METIS achieves $1.64-2.54\times$ lower response delay compared to fixed configurations of comparable quality.
  • $\bullet$
    Higher throughput : METIS achieves $1.8-4.5\times$ higher throughput than RAG serving systems which use fixed configurations reaching similar quality.
  • $\bullet$
    Negligible overhead : METIS’ profiler’s delay is negligible compared to the overall delay of the LLM’s RAG inference.
Figure 10. METIS achieves $1.64-2.54\times$ lower delay compared to both best fixed configuration baselines and quality-optimized RAG configuration without sacrificing generation quality.
Figure 11. METIS achieves $1.8-4.5\times$ higher throughput (at 1.8 seconds) than baselines which use fixed configurations of closest (not higher) quality.
7.1. Setup
Models and hardware: : In RAG, models use external context to answer queries instead of the trained weights (model’s embedded knowledge). The model extracts data from the external context and has stringent serving latency requirements. Hence RAG applications use smaller, instruction-tuned models. We evaluate METIS on a popular models for RAG LLM inference, specifically the fine-tuned version of Mistral-7B-v3 and Llama3.1-70B for additional experiments as they are commonly used in all RAG QA workload serving. All models are fine-tuned such that they can take long contexts (up to 32K and 128K respectively). We apply AWQ-model quantization to both models.
We use an NVIDIA A40 GPU server with 2 GPUs to benchmark our results. The server is equipped with 384GB of memory and two Intel(R) Xeon(R) Gold 6130 CPUs with Hyper-threading and Turbo Boost enabled by default. We use 1 GPU to serve Mistral-7B-v3 and 2 GPUs to serve Llama3.1-70B.
Datasets: We use multiple RAG QA datasets with various query profiles, in order to have task-representative workloads. Table 1 summarizes their input-output statistics.
  • $\bullet$
    Squad (squad, ): Squad is a single-hop reading comprehension dataset, consisting of questions on articles, where the answer to every question is a segment from the corresponding reading passage.
  • $\bullet$
    Musique (musique, ): Musique is a multihop QA dataset with reasoning-based questions. It is designated to test LLM’s reasoning ability where one reasoning step critically relies on information from another.
  • $\bullet$
    KG RAG FinSec (kgrag, ): KG RAG Finsec is part of a Knowledge Graph family of RAG datasets and focuses on financial domain questions from Fortune 500 companies. This dataset contains quarterly financial reports and queries need to read information for multiple chunks for answering.
  • $\bullet$
    QMSUM (qmsum, ): QMSUM is a human-annotated query-based multi-domain meeting summarization benchmark designed to test LLM’s reasoning-based summarization capabilities. This dataset contains multiple meeting transcripts and queries to summarize relevant spans of meetings.
We build a retrieval database database by splitting the queries’ contexts into fixed-sized chunks using Langchain (Chase_LangChain_2022, ) for the database, with Cohere embed-v3.0 (cohere_embed, ) embeddings and FAISS (douze2024faiss, ) L2-distance similarity search in order to retrieve relevant chunks for RAG inference. To simulate a real RAG workload, we choose 200 queries from each dataset, and send them concurrently to METIS using a Poisson distribution with an average arrival rate of 2 per dataset . We report the results per dataset.
@@ 7.1. Setup
Table 1. Input and output length (# of tokens) distributions of the RAG datasets used in our evaluation.
@@
Figure 12. Understanding the delay improvement in METIS
Baselines: We compare METIS with the following baselines.
  • vLLM: We serve RAG with vLLM with multiple static configurations across different queries.
@@
  • Parrot*: We implement Parrot’s (lin2024parrotefficientservingllmbased, ) configuration-based batching. Parrot* does not adapt the configuration per query. We compare with Parrot* using fixed RAG configurations which achieve the closest quality to us.
  • AdaptiveRAG*: We implement AdaptiveRAG’s (jeong-etal-2024-adaptive, ), query complexity-based RAG-configuration selection and choose the configuration which maximizes the F1-score, without considering the system resource cost.
7.2. Overall improvement
@@
Lower delay without sacrificing generation quality: Figure 10 shows METIS achieves delay reduction $1.64-2.54\times$ over AdaptiveRAG* with no reduction in F1-score. Over using fixed configurations of similar delay, served with both Parrot* and vLLM, METIS achieves $12-18\%$ higher F1-score.
@@
Higher throughput at lower delay: Figure 11 shows METIS achieves higher throughput compared to fixed configuration baselines when they choose the fixed-config which achieves the closest quality. METIS has dynamic RAG configurations while vLLM and Parrot only allow a fixed configuration across queries. Baselines choose the configuration which achieves the highest average F1-score among all the fixed configurations but due to the static nature, every configuration achieves a lower F1-score compared to METIS Compared to Parrot* and vLLM, METIS achieves $1.8-4.5\times$ times higher throughput.
@@
Understanding METIS’ improvement: METIS’s gains come from jointly selecting the configuration based on the available resource, along with performing scheduling. METIS achieves higher quality than the fixed-config baselines as it is adapts the RAG-configuration per query. It reduces delay by resource-aware scheduling, making it better than fixed configurations which achieve closest quality.
@@
METIS achieves higher throughput as it is able to adapt configurations based on resource availability as compared to the baselines. Both Parrot* and vLLM schedule fixed RAG-configurations and cannot benefit from delay achieved by adapting the configuration like METIS. Parrot* can improve the delay over using fixed configurations with vLLM by $1.4-1.8\times$ but cannot improve the quality. Compared to AdaptiveRAG*, METIS achieves lower latency, as AdaptiveRAG* inflates serving latency without considering the cost of profiling or the cost of the configuration. AdaptiveRAG* also does not provide an interface to extend to multiple knobs.
7.3. Analyzing the gains from METIS
@@
Delay saving: Figure 12 shows the contribution of every component of METIS. We compare with vLLM’s fixed configuration, which achieves the highest quality (blue bar). Using the profiler’s outputs and choosing the median value every time (orange bar), we achieve $1.4-1.68\times$ reduction in delay. Next, we see the effect of batching (like Parrot*), by choosing the median value configuration and batching, we achieve $1.1-1.2\times$ reduction in delay. Finally, METIS achieves even greater delay reduction by $1.45-1.75\times$ by adapting the configuration based on available GPU memory with batching.
@@
Figure 13. Even with increasing the inference model size, fixed configurations have $2.38-6.8\times$ higher cost and lower quality compared to METIS.
@@
Figure 14. Improvement for METIS using feedback from the output helps improve the F1-score by $4-6\%$.
@@
Figure 15. METIS achieves lower delay by $2.1-2.4\times$ at the same quality even with a larger inference LLM.
@@
Figure 16. Breakdown analysis: By tuning more knobs in METIS, we can see better quality-delay tradeoffs.
@@
Figure 17. METIS’ performance gains remain substantial even with a smaller, open-source LLM profiler.
@@
Figure 18. METIS’ profiler delay is at most 1/10th of end-to-end response delay across queries from all datasets.
@@
Cost saving: Figure 13 shows METIS (including its profiler) has significant lower dollar cost and higher F1-score, compared to choosing the best fixed configuration, with increasing model complexity. The cost of using a (LLama3-70B) inference model with vLLM and a fixed configuration is higher by $2.38\times$ times while also having a lower F1-score of $6.5\%$ times across datasets. Even more powerful inference models like GPT-4o fail to achieve the same F1-score with fixed configurations but have a much higher cost of $6.8\times$.
@@
Profiler feedback-based improvement: In Figure 14 we show the effect of the golden-configuration-based feedback to the profiler in order to improve its output. We use a 350 query sample for the QMSUM and KG RAG FinSec dataset as the workload. We see that with the feedback mechanism (blue line), the F1-score improves by $4-6\%$ as compared to not having feedback (red line) from the outputs of the golden configuration. We ensure that the feedback mechanism cannot result in the output of very expensive configurations, as METIS’ joint scheduler will not pick increasingly expensive configurations based on the GPU resource constraint.
7.4. Sensitivity analysis
@@
Changing the inference LLM: Figure 15 shows the outcome of changing the inference LLM to a larger LLM (Llama3.1-70B) on the Musique and QMSUM datasets. Even with a more powerful LLM, METIS achieves $2.1-2.4\times$ lower delay than AdaptiveRAG* at a similar F1-score. The best fixed-configuration baselines such as Parrot* and vLLM have a lower F1-score of $7-10\%$. In RAG, models mainly rely on the external context to answer the question instead of the model weights and we only get a $2\%$ improvement in F1-score compared to the smaller inference models.
@@
Incrementally tuning knobs in METIS: In Figure 16, we show the benefit we the improvement we get by incrementally adding more knobs to METIS. We measure this for the QMSUM dataset with the original Mistral-7B-v3 model. We first only tune the num_chunks (red point). Progressively we tune the RAG-configuration knobs of synthesis_method and intermediate_length and scheduling. We achieve $5,4,3\%$ higher F1-Score compared to vLLM. Finally, by adding the scheduling, $2.8\times$ lower delay reduction in delay.
@@
Changing the profiler LLM: Figure 17 shows the effect of changing the LLM profiler from GPT-4o to a smaller Llama3.1-70B model. METIS with the new profiler, still achieves $1.4-2.1\times$ over AdaptiveRAG* with a similar F1-score. Static configurations of Parrot* and vLLM which achieve similar delay, METIS achieves $10-14\%$ higher F1-score.
@@
Delay overhead of METIS’s per-query profiling: We show the negligible delay overhead of using an LLM profiler within METIS. Figure 18 shows the fraction of METIS’ profiler of the total end-to-end delay. Using the profiler at most adds $0.1$ fraction and in the average case only adds 0.030.060.03-0.06 fraction to the total delay across queries from all datasets.
Figure 19. METIS’ under low load without batching
METIS’ performance under low load: In Figure 19, we evaluate METIS under low load by sending queries sequentially, with every query sent after the previous query is completed. We compare with vLLM’s fixed configuration, which achieves the highest quality (blue bar). METIS uses its best-fit algorithm to pick the most expensive configuration from the pruned space of configurations. As METIS only picks from configurations relevant to the query profile, it is still able to reduce delay by $1.48-1.56\times$ under low load.
8. Related work
@@
Systems for serving RAG: Several systems have been proposed for RAG (lin2024parrotefficientservingllmbased, ; jin2024ragcacheefficientknowledgecaching, ; tan2024teolaendtoendoptimizationllmbased, ; martin2024llmproxyreducingcostaccess, ; kim2024effectschedulingpreemptionefficiency, ; jeong-etal-2024-adaptive, ; zhang2024pqcacheproductquantizationbasedkvcache, ; edge2024localglobalgraphrag, ; infercept, ; piperag, ; ralmspec, ) which focus on improving retrieval using complex, iterative retrieval algorithms or on serving model selection. METIS can work in conjunction with such systems as METIS focuses on optimizing quality and serving latency, independent of how the retrieval algorithm identifies chunks for retrieval.
@@
KV cache storage and retrieval: Storing and reusing KV cache across different requests have been commonly studied in recent work (10.1145/3651890.3672274, ; hu2024memservecontextcachingdisaggregated, ; liu2024minicachekvcachecompression, ; dong2024lesssynthesizingrecurrencekv, ; liu2024droidspeakenhancingcrossllmcommunication, ; pan2024marconi, ; infercept, ; cachedattention, ; pensieve, ; preble, ; zheng2024sglang, ; kwon2023efficient, ; jiang2025ragosystematicperformanceoptimization, ). METIS can work alongside these systems, where instead of retrieving chunks, it can retrieve the KV Caches for generating the output. In RAG, some additional optimizations are needed to combine KV Caches of different chunks that don’t share a common prefix. This is important as the trivial concatenation of KV Caches loses important cross-attention and reasoning between chunks. These optimizations are enabled by KV Cache blending-based approaches (yao2024cacheblendfastlargelanguage, ; hu2024epicefficientpositionindependentcontext, ; cheng2024largelanguagemodelsneed, ; jin2024computeloadkvcache, ; wang2024modeltellsmergeadaptive, ; promptcache, ). However RAG workloads have a large number of related contexts across queries and storing all the KV Cache is extremely expensive. We do not measure the KV Cache reuse ratio across queries and leave it for future work.
@@
Prefill-Decode Optimizations: Several systems have proposed optimizations to speed-up prefill and decode for LLMs by leveraging unique properties of each phase (sarathi, ; distserve, ; 10.1145/3694715.3695948, ; qin2024mooncakekvcachecentricdisaggregatedarchitecture, ; jiang2024neosavinggpumemory, ; apparate, ; powerinfer, ; nanoflow, ). Notable techniques include chunked-prefill which allows interleaving prefill and decode requests and disaggregated prefill which separates compute nodes for prefill and decode. All of these optimizations enable faster generation speed but don’t focus on generation quality. METIS can be applied with such LLM serving systems optimizations.
9. Discussion and Limitations
While METIS’ profiler and configuration mapping algorithm is currently designed to work with commonly deployed RAG QA pipelines, it can be easily extended to generalize across new RAG configurations, workflows and domains.
Agentic RAG: New research directions in RAG (zhang2024raftadaptinglanguagemodel, ; li2025agenticragdeepreasoning, ; guo2025reagannodeasagentreasoninggraphagentic, ) have developed pipelines with LLM agents, tool calling and deep chain-of-thought to be used for RAG workloads. For an agentic workflow, a key extension for METIS is to profile the query-complexity and break down a query into multiple sub-queries for planning (e.g., how many sub-queries are needed becomes a new configuration knob). METIS complements such workflows and can continue to perform the joint resource allocation for each sub-query.
Multi-modal RAG: Emerging multi-modal LLMs has led to the need for multi-modal RAG (riedler2024textoptimizingragmultimodal, ; liu2025hmraghierarchicalmultiagentmultimodal, ; hu2025mragbenchvisioncentricevaluationretrievalaugmented, ) and METIS can be extended to complement this. Using the natural-language profile of the query, METIS can profile the different types of data to be retrieved. e.g., A query might ask for a fact and a supporting image and this becomes a new configuration knob. Based on the properties of the new data retrieved, a configuration selection mapping rule can be added to decide which final RAG configuration should be chosen.
Graph-based RAG: Another emerging area is GraphRAG (10.1145/3711896.3737012, ; edge2024localglobalgraphrag, ; zhou2025indepthanalysisgraphbasedrag, ) where retrieval is performed by choosing from hierarchical communities (e.g., coarse grained data aggregation vs fine-grained facts). METIS can complement such approaches by using the complexity profile of the query it generates, in order to choose the appropriate depth of community to use, and add this as a configuration knob.
10. Conclusion